# Artificial Neural Networks for Magnetoencephalography: A review of an emerging field

### v1

@@ 4.3 Current limitations and challenges
The diversity and wide range of ways ANNs have been applied to MEG research implies that the current challenges and limitations are also very diverse. Nevertheless, we will now attempt to categorize the dominant issues we observed across the reviewed papers into a handful of topics. Whenever possible, we offer strategic guidance to navigate these challenges in the hope of enhancing the robustness, efficacy, and reproducibility of future research in this rapidly evolving field. The wide range of ways ANNs have been applied to MEG has revealed a panorama of challenges and limitations as diverse as the methodologies themselves. In an attempt to synthesize for the reader the main challenges we identified across the reviewed literature, we categorize the predominant issues and pitfalls into a concise set of topics. First, we will address the general data scarcity problem prevalent in MEG-based neuroimaging studies. Next, we will underscore the importance of model validation techniques and the need for an accurate performance evaluation. We will then discuss the importance of hyperparameters in ANN-based MEG studies. Lastly, we will discuss the issue of reproducibility in ANN-based MEG studies.
@@ 4.3.1 Dealing with data scarcity
It is recommended that data augmentation techniques be used when feasible to cope with scarce data. For event-based epochs, introducing different event timings for each trial window will increase the sample size. In the case of continuous data, it is possible to use overlapping segments to artificially increase dataset size. Depending on the type of study and data, various data augmentation techniques can be added to a study’s preprocessing pipeline. For more details on the available methods, we refer the reader to [152], in which the techniques used for EEG data augmentation can also be applied to MEG data.
@@ 4.3.2 Model validation techniques
In many neural network applications, the vast amount of data combined with the typical training duration for ANNs makes cross-validation an impractical method for model validation. Instead, most of the time, the dataset is randomly split into three subsets to validate the model: training, validation, and test (also called evaluation set). The training set generally represents 50 to 80 percent of the initial dataset, and the remaining data is split in half to create the validation and test sets. The training set is employed to train the network with a set of hyperparameters and tested on the validation set. The test set is only used at the end when the architecture and hyper-parameters are fixed and once the desired performance has been reached on the validation set. It is solely used to evaluate the final model performance. However, when data is scarce, as it is often the case for neuroscience studies, it is recommended to use re-sampling methods such as bootstrapping or cross-validation to make use of the data as much as possible [138]. Cross-validation is often preferred (see Varoquaux’s response to [138]). It will allow the train and validation sets to be changed several times to acquire a less noisy measure of the model’s performance. In the studies included in this survey, when cross-validation is used, K-fold was the preferred technique, with the value of K ranging from 4 to 50. When the goal is to test generalization to new subjects, the "Leave P subjects out" cross-validation or a stratified version of K-Fold is recommended.
@@ 4.3.3 Model performance evaluation and baseline comparisons
Some classification studies did not include baseline results against which one could have compared the reported ANN performances. This can be a relevant concern given that classical shallow ML approaches (e.g. logistic regression ([153]), Random forest [154]), SVM ([155]), etc.) are generally easier and faster to implement than ANNs. Generally, one should always evaluate simple solutions before spending time and effort on more complex ones, which might also be more challenging to interpret. Similarly, only a few articles among the modeling studies mentioned some kind of baseline performance comparison. Ideally, modeling studies should report Noise Ceiling (NC) and/or evaluate the performance of their similarity analysis with an untrained model. Together, these measures would reinforce robust similarity scores’ reliability and help identify and discard weaker results.
@@ 4.3.4 Hyperparameter settings
In recent years, we have witnessed the development of increasingly deep networks aimed, for a large part, at enhancing performance in image classification tasks. However, the studies reviewed here reveal a trend toward relatively shallower network architectures. Intriguingly, when depth is evaluated as a hyperparameter in these studies, it appears to have a lesser impact on performance compared to other parameters, such as layer width or filter size. When using ANNs, the training phase is a pivotal component, and an incorrect selection of the optimizer and parameters could result in a network that fails to learn. Unfortunately, 16 of the classification studies did not provide enough information about the hyperparameters used. However, most studies specify the optimizer, learning rates, batch size, and the number of epochs, including the early stop condition when applicable. It is essential to systematically detail a model’s hyperparameters and training regimen (in any type of ANN study) to ensure the ability to reproduce results or to use the same model similarly with different data. In this context, a recommended practice is to also provide the source code alongside all the parameters and hyperparameters used to generate any given result of a study. Lastly, it’s worth mentioning that many studies in our review employ batch sizes that are not powers of two. However, using batch sizes that are powers of two is recommended to optimize computation and reduce training times [156].
@@ 4.3.5 The issue of reproducibility
In the field of neuroscience, concerns about the reproducibility of results are prevalent. The substantial costs and privacy issues related to acquiring neuroimaging or behavioral data restrict the availability of open-access datasets, thereby challenging the reproducibility of research conducted with such data ([157, 158]). In addition, this challenge is compounded by numerous studies that did not include enough information about the network architecture or the training parameters or did not provide access to their code, from which this information could be extracted. This lack of transparency further impedes the reproducibility of the results, making it difficult for researchers to validate and build on previous findings. The application of ML to neuroscience data faces challenges, mainly due to the relatively small/moderate sample sizes in neuroscience datasets. Besides, most algorithms require large amounts of data to achieve robust performance. Moreover, small datasets will make generalization harder to reach since having fewer data will further increase the impact of inter-subject variability. Solutions exist to tackle problems caused by small datasets, such as cross-validation, bootstrapping, and data augmentation. Still, more data will always be a preferable approach to achieve better results ([1]). However, the trend of small private datasets in the world of neuroscience is slowly changing with increasing awareness around the importance of open science practices, and public datasets are becoming more common.
@@ 4.3.6 Interpretation and visualizations
In Brain-Computer Interface (BCI) studies, where the primary concern is the algorithm’s performance, the black-box nature of ANNs is often considered less problematic. However, in many other applications, including brain decoding studies, gaining insight into what is happening in the model’s latent space is key to understanding how the information was extracted from the signal, as well as why and how decoding is possible. In domains like image classification or object detection, there are various tools (Grad-CAM, tSNE ([159]), UMAP ([160]) that allow the user to have a representation of what the network ’sees’ and what properties of the data are important in its decision-making process. A large proportion (52 out of 69) of the classification studies, where interpretation would be the most interesting, did not use these tools to gain insights into the representations learned and used by the model.
@@ 4.3.7 Note of caution on inconsistent nomenclature
Consistent terminology is, of course, vital for effective communication across all scientific disciplines. However, it becomes particularly crucial in emerging fields where the community still defines methodologies and core concepts. Adopting a uniform nomenclature is essential in a rapidly evolving research domain, such as the intersection of ANN and neuroimaging. Our extensive literature review reveals that terminology within classification studies is fairly consistent. However, we noted that the nomenclature in other categories, particularly within modeling studies, tends to show less consistency. To give a few examples, some research papers include RSA as a component of the MVPA analysis, while others do not. Additionally, some studies use neural predictability but do not explicitly mention it in their papers. In the modeling studies reviewed, the analysis pipelines exhibit considerable variety and flexibility, leading to notable diversity across the studies. For researchers looking to adopt these methodologies, we recommend a thorough examination of multiple studies. This approach will provide a comprehensive understanding of the various pipelines and help identify the methods most aligned with specific research objectives. Additionally, it is crucial to strive for consistency in terminology with the existing literature to ensure clarity and help the community move towards established procedures and nomenclature. For those interested in RSA in particular, insightful comments and critiques of the method can be found in [161].
@@ 5 Conclusion
Machine learning, notably the use of artificial neural networks, has become a cornerstone in contemporary research, reshaping both data analysis and modeling across diverse scientific fields. Although ANNs are relatively novel in the realm of neuroimaging, the collection of studies in this review demonstrates a rapidly growing interest in employing ANNs for both the analysis and modeling of MEG data. This trend is part of a broader movement towards integrating advanced data-driven and computational methods to deepen our understanding of neural mechanisms. The large body of work we reviewed indicates that ANNs are increasingly recognized as powerful tools in neuromagnetic imaging due to their robust performance, versatility, and ability to handle complex datasets. A comparison between our figure 1 and Figure 10b in [17] suggests a trajectory similar to that observed in the field of EEG, where ANNs have already shown extensive utility. This parallel points to a likely expansion in the number of studies employing ANNs for MEG data, further diversifying the applications and methodologies within the field.
One of the most remarkable observations is the incredibly diverse ways in which ANNs have so far been applied to MEG. While shallow ML tools like SVM or random forest classifiers are mainly applied to MEG in the context of brain decoding studies, ANNs open up a wide range of applications. As shown here, these extend from addressing the source estimation problem to providing new tools for MEG data cleaning and artifact rejection. Furthermore, our survey indicates an increasing interest within the neuroscience community in employing ANNs to build and evaluate information processing models in the human brain, reflecting the expanding role of these networks in advancing neuroscientific research [162].
@@
Despite the remarkable advances in AI tools and the new opportunities that Artificial Neural Networks (ANNs) present for MEG data analysis and modeling, it is critical to avoid the allure of AI hype and the rush to employ complex AI algorithms solely for their novelty. Many research questions may still be best addressed using standard approaches that are well-established within the field and which often come with greater result interpretability. Moreover, the extensive range of parameters and hyperparameters involved in ANNs, coupled with the expertise required for their proper implementation, can sometimes introduce errors or yield misleading outcomes.
Regarding future developments, it seems that many exciting avenues are unfolding. To name just two of these: First, we believe that advances in AI interpretability and visualization tools will be key to enhancing the utility of ANNs in MEG research. Second, the recent trend towards AI foundation models in many disciplines, including neuroscience, will certainly also play an increasingly helpful role in MEG, both for basic and clinical applications. The scientific community is showing a growing interest in developing foundation models for brain imaging modalities ([163, 164]) and time series data ([165]). We anticipate that the neuroscience community will follow this trend ([121, 84]). Crucially, foundation models provide a convenient framework to combine multimodal data. Time will tell; one of the next paradigm shifts in this field will come from combining imaging modalities like MEG, EEG, and fMRI into a single multi-purpose foundation model.
@@
Looking ahead, the future of this emerging research appears ripe with promising developments. Two particularly exciting prospects stand out. First, advances in AI interpretability and visualization tools are poised to significantly enhance the application of ANNs in MEG research, making these complex models more accessible and their findings more actionable. Second, the burgeoning adoption of AI foundation models across various disciplines, including neuroscience, is set to play a transformative role in both basic and clinical MEG applications. The scientific community’s growing enthusiasm for developing foundational models for brain imaging modalities ([163, 164]) and time series data ([165]) suggests that neuroscience will not be far behind ([121, 84]). These foundation models offer a versatile framework for integrating multimodal data, paving the way for a potential paradigm shift that could merge imaging modalities like MEG, EEG, and fMRI into a unified, multipurpose model. Integrating such modalities could revolutionize our understanding of the brain and its disorders.

### v3

@@ 4.3 Current limitations and challenges
The diversity and wide range of ways ANNs have been applied to MEG research reflect a correspondingly broad spectrum of challenges and limitations. To provide clarity, we categorize the predominant issues observed across the reviewed literature into a concise set of topics, offering strategic guidance where possible to enhance the robustness, efficacy, and reproducibility of future work in this rapidly evolving field. First, we address the general data scarcity problem prevalent in MEG-based neuroimaging studies. Next, we underscore the importance of model validation techniques and the need for accurate performance evaluation. We then discuss the critical role of hyperparameters in ANN-based MEG studies. Lastly, we examine the issue of reproducibility in this domain.
@@ 4.3.1 Dealing with data scarcity
It is recommended that data augmentation techniques be used when feasible to cope with scarce data. For instance, common approaches for time-series data like M/EEG include adding noise, applying transformations such as amplitude scaling or time warping, or generating synthetic trials using generative models. For event-based epochs, introducing different event timings for each trial window will increase the sample size. In the case of continuous data, it is possible to use overlapping segments to artificially increase dataset size. Depending on the type of study and data, various data augmentation techniques can be added to a study’s preprocessing pipeline. For a systematic comparison and more details on available methods specifically evaluated on EEG, many of which are applicable to event-related MEG data, we refer the reader to [200].
@@ 4.3.2 Model validation techniques
In many neural network applications, the vast amount of data combined with the typical training duration for ANNs makes cross-validation an impractical method for model validation. Instead, most of the time, the dataset is randomly split into three subsets to validate the model: training, validation, and test (also called evaluation set). The training set generally represents 50 to 80 percent of the initial dataset, and the remaining data is split in half to create the validation and test sets. The training set is employed to train the network with a set of hyperparameters and tested on the validation set. The test set is only used at the end when the architecture and hyper-parameters are fixed and once the desired performance has been reached on the validation set. It is solely used to evaluate the final model performance. However, when data is scarce, as it is often the case for neuroscience studies, it is recommended to use re-sampling methods such as bootstrapping or cross-validation to make use of the data as much as possible ([177]). It will allow the train and validation sets to be changed several times to acquire a less noisy measure of the model’s performance. In the studies included in this survey, when cross-validation is used, K-fold was the preferred technique, with the value of K ranging from 4 to 50. When the goal is to test generalization to new subjects, the "Leave P subjects out" cross-validation or a stratified version of K-Fold is recommended.
@@ 4.3.3 Model performance evaluation and baseline comparisons
Some classification studies did not include baseline results against which one could have compared the reported ANN performances. This can be a relevant concern given that classical shallow ML approaches (e.g. logistic regression [201], random forest [179], SVM [202], etc.) often involve simpler classifier implementations and potentially faster training times —once appropriate features have been engineered—, compared to designing and training complex ANNs. However, the overall practical advantage depends heavily on the computational complexity of the necessary feature extraction pipeline, which can itself be intensive and time-consuming, potentially offsetting the gains in classifier simplicity. Additionally, it is critical to evaluate classification performance against chance-level baselines, particularly in studies with small sample sizes, where theoretical chance levels (e.g., 50% accuracy for two-class problems) can be misleadingly optimistic. Limited data can lead to inflated accuracies, even exceeding 70%, purely by chance and rigorous statistical evaluations, such as permutation tests, are required to properly establish meaningful baseline performances, thus ensuring robust assessments of classifier significance [203]. Generally, one should always evaluate simple solutions before spending time and effort on more complex ones, which might also be more challenging to interpret. Similarly, only a few articles among the modeling studies mentioned some kind of baseline performance comparison ([121, 122, 118, 119, 126, 129]). Ideally, modeling studies should report NC and/or evaluate the performance of their similarity analysis with an untrained model. Together, these measures would reinforce robust similarity scores’ reliability and help identify and discard weaker results.
@@ 4.3.4 Hyperparameter settings
In recent years, we have witnessed the development of increasingly deep networks aimed, for a large part, at enhancing performance in image classification tasks. However, the studies reviewed here reveal a trend toward relatively shallower network architectures. Intriguingly, when depth is evaluated as a hyperparameter in these studies, it appears to have a lesser impact on performance compared to other parameters, such as layer width or filter size. When using ANNs, the training phase is a pivotal component, and an incorrect selection of the optimizer and parameters could result in a network that fails to learn. Unfortunately, 16 of the classification studies did not provide enough information about the hyperparameters used. However, most studies specify the loss function, optimizer, learning rates, batch size, and the number of epochs, including the early stop condition when applicable. It is essential to systematically detail a model’s hyperparameters and training regimen (in any type of ANN study) to ensure the ability to reproduce results or to use the same model similarly with different data. In this context, a recommended practice is to also provide the source code alongside all the parameters and hyperparameters used to generate any given result of a study. Furthermore, given the dependence of optimal hyperparameters, including network depth, on the specific application, researchers designing networks for their own MEG studies may find the detailed information collated in tables 2, 3, 4, and 5 useful as a starting point or source of inspiration. Lastly, it’s worth mentioning that many studies in our review employ batch sizes that are not powers of two. However, using batch sizes that are powers of two is recommended to optimize computation and reduce training times ([204]).
4.3.5 Variability in MEG data acquisition, preprocessing, and experimental protocols
One of the main challenges in applying artificial neural networks (ANNs) to MEG research is the variability in acquisition hardware, preprocessing steps, and experimental protocols across studies. These differences can affect data quality and consistency, which in turn influence model performance and the ability to compare results across studies.
At the acquisition level, studies used different sensor types (e.g., magnetometers, planar gradiometers, axial gradiometers) and MEG systems (e.g., Elekta Neuromag, CTF, BTi), with sampling frequencies ranging from 50 Hz to 2400 Hz. Preprocessing methods also varied: some studies used ICA (often via EEGLAB or FieldTrip) for artifact rejection, others applied wavelet denoising or relied on proprietary tools. In 39 studies, preprocessing steps were not clearly reported. Segment lengths and filtering parameters also differed, even in studies using similar tasks.
Experimental protocols varied in terms of task design (e.g., resting-state, oddball, language processing, or BCI tasks), number and timing of trials, baseline usage, and whether the data were epoched or treated as continuous. This variability makes it difficult to directly compare ANN architectures or performance metrics.
To address these challenges, future studies would benefit from more consistent preprocessing pipelines, standardized task protocols where feasible, and clearer reporting of methods. Efforts like BIDS-MEG and open datasets with thorough documentation can support better reproducibility and enable more reliable benchmarking across studies.
@@ 4.3.6 The issue of reproducibility
In the field of neuroscience, concerns about the reproducibility of results are prevalent. The substantial costs and privacy issues related to acquiring neuroimaging or behavioral data restrict the availability of open-access datasets, thereby challenging the reproducibility of research conducted with such data ([205, 206]). In addition, this challenge is compounded by numerous studies that did not include enough information about the network architecture or the training parameters or did not provide truly open access to their code, from which this information could be extracted. While some studies state code is ’available upon request’, this often proves ineffective in practice, with requests frequently going unanswered or being declined ([207]), meaning truly accessible code is likely rarer than implied. This overall lack of transparency and accessibility further impedes the reproducibility of the results, making it difficult for researchers to validate and build on previous findings. The application of ML to neuroscience data faces challenges, mainly due to the relatively small/moderate sample sizes in neuroscience datasets. Besides, most algorithms require large amounts of data to achieve robust performance. Moreover, small datasets will make generalization harder to reach since having fewer data will further increase the impact of inter-subject variability. Solutions exist to tackle problems caused by small datasets, such as cross-validation, bootstrapping, and data augmentation. Still, more data will always be a preferable approach to achieve better results ([1]). However, the trend of small private datasets in the world of neuroscience is slowly changing with increasing awareness around the importance of open science practices, and public datasets are becoming more common.
The utility of these public datasets for reproducible research and cross-study comparisons is greatly enhanced by standardized data formats and organizational structures, notably the Brain Imaging Data Structure extension for MEG (MEG-BIDS) [208]. Resources providing standardized, open-access MEG recordings suitable for ANN analysis include repositories and datasets such as OpenNeuro (https://openneuro.org/), the human connectome project (https://www.humanconnectome.org/), the Cambridge centre for ageing and neuroscience (Cam-CAN) dataset (https://camcan-archive.mrc-cbu.cam.ac.uk/dataaccess/), the open MEG archive (OMEGA) database (https://www.mcgill.ca/bic/neuroinformatics/omega), and the MNE sample datasets (https://mne.tools/stable/overview/datasets_index.html), offering valuable benchmarks for training, testing, and comparing models.
@@ 4.3.7 Interpretation and visualizations
In BCI studies, where the primary concern is the algorithm’s performance, the black-box nature of ANNs is often considered less problematic. However, in many other applications, including brain decoding studies, gaining insight into what is happening in the model’s latent space is key to understanding how the information was extracted from the signal, as well as why and how decoding is possible. In domains like image classification or object detection, there are various tools (Grad-CAM [209], tSNE [210], UMAP [211]) that allow the user to have a representation of what the network ’sees’ and what properties of the data are important in its decision-making process. A large proportion (52 out of 70) of the classification studies, where interpretation would be the most interesting, did not use these tools to gain insights into the representations learned and used by the model. A similar pattern of limited application was observed in the ’Other’ category studies reviewed. While interpretability might be considered less critical for some preprocessing or source localization tasks compared to decoding, it remains vital for evaluating novel methodological approaches. Indeed, roughly half of the ’Methods’ studies (9 out of 16) included visualization techniques such as filter or activation map analysis. Interpretability tools were also used in a majority of the ’Preprocessing’ studies (5 out of 8), often employing methods like Grad-CAM [182]. However, interpretability techniques were notably absent in all reviewed ANN-based source localization studies. This suggests that developing and applying robust interpretability methods beyond classification and basic modeling tasks remains an important area for future work in MEG research using ANNs.
@@ 4.3.8 Note of caution on inconsistent nomenclature
Consistent terminology is important for effective communication across all scientific disciplines. However, it becomes particularly crucial in emerging fields where the community still defines methodologies and core concepts. Adopting a uniform nomenclature is essential in a rapidly evolving research domain, such as the intersection of ANN and neuroimaging. Our extensive literature review reveals that terminology within classification studies is fairly consistent. However, we noted that the nomenclature in other categories, particularly within modeling studies, tends to show less consistency. To give a few examples, some research papers include RSA as a component of the MVPA analysis, while others do not. Additionally, some studies use neural predictability but do not explicitly mention it in their papers. In the modeling studies reviewed, the analysis pipelines exhibit considerable variety and flexibility, leading to notable diversity across the studies. For researchers looking to adopt these methodologies, we recommend a thorough examination of multiple studies. This approach will provide a comprehensive understanding of the various pipelines and help identify the methods most aligned with specific research objectives. Additionally, it is crucial to strive for consistency in terminology with the existing literature to ensure clarity and help the community move towards established procedures and nomenclature. For those interested in RSA in particular, insightful comments and critiques of the method can be found in [212].
4.3.9 Hardware and practical application constraints
Another consideration involves the practical constraints of MEG for certain applications, notably BCIs. While EEG is often considered more practical for widespread BCI use due to lower cost and portability, MEG’s arguably superior signal quality and spatial resolution offer potential advantages for specific BCI paradigms requiring high fidelity decoding, as explored in several studies reviewed here ([97, 99, 95]). However, traditional SQUID-based MEG systems face significant hurdles for practical BCI deployment, including high cost, the need for magnetically shielded rooms, and strict subject immobility requirements. Although ANNs can contribute by improving the robustness of decoding complex MEG signals and potentially adapting to noise or variability (3.2), they do not resolve these fundamental infrastructural and hardware limitations. Promisingly, emerging sensor technologies like optically pumped magnetometers (OPMs) may mitigate these physical constraints. OPMs operate without cryogenics, are wearable, and allow for subject movement, potentially enabling more flexible, cost-effective, and practical MEG-based BCI systems in the future ([213]).
4.3.10 Challenges in ANN-based source reconstruction
Applying ANNs to the MEG inverse problem, i.e., source reconstruction, presents both opportunities and significant challenges, as highlighted by the studies reviewed in section 3.4. Source reconstruction is inherently ill-posed and highly sensitive to factors like sensor noise, assumptions about source activity, and, crucially, the accuracy of the head model used to compute the forward solution ([8, 29]). While ANNs offer potential advantages, such as learning complex non-linear source-sensor mappings directly from data, their effectiveness, particularly when trained on simulated data as is common practice ([167, 137]), is fundamentally constrained by the fidelity of these simulations.
Indeed, if an ANN is trained using data generated with an inaccurate forward model (due to simplified head geometry, incorrect conductivity values, or sensor misalignment), its ability to generalize and accurately localize sources in real-world MEG data will be severely compromised. The network may perform well on simulated test data derived from the same flawed model but fail when applied to actual measurements. Therefore, the development of effective ANN-based source reconstruction methods heavily relies on the generation of high-quality training data using realistic simulations and, most importantly, accurate subject-specific forward models. Furthermore, even with accurate models, challenges remain in ensuring that ANNs do not simply learn biases present in the simulation protocols and in interpreting the learned source representations. It is therefore important to keep in mind that although ANNs may provide a powerful data-driven alternative to traditional inverse methods, they do not circumvent the fundamental physical constraints and modeling requirements inherent to MEG source localization.
@@ 5 Conclusion
Machine learning, notably the use of ANNs, has become a cornerstone in contemporary research, reshaping both data analysis and modeling across diverse scientific fields. Although ANNs are relatively novel in the realm of neuroimaging, the collection of studies in this review demonstrates a rapidly growing interest in employing ANNs for both the analysis and modeling of MEG data. This trend is part of a broader movement towards integrating advanced data-driven and computational methods to deepen our understanding of neural mechanisms. The large body of work we reviewed indicates that ANNs are increasingly recognized as powerful tools in neuromagnetic imaging due to their robust performance, versatility, and ability to handle complex datasets. A comparison between our figure 2 and figure 10b in [5] suggests a trajectory similar to that observed in the field of EEG, where ANNs have already shown extensive utility. This parallel points to a likely expansion in the number of studies employing ANNs for MEG data, further diversifying the applications and methodologies within the field.
One of the most remarkable observations is the diverse range of applications for which ANNs are now being employed with MEG data. While traditional ML tools like SVM or random forest classifiers are well-established, particularly for brain decoding and classification tasks based on extracted features ([48]), ANNs facilitate a distinct spectrum of applications, as demonstrated throughout this review. Notably, their capacity for representation learning and building end-to-end systems has led to their use not only in classification but also in addressing the source estimation problem ([167, 137]), providing novel tools for data cleaning and artifact rejection ([147, 148]), and, significantly, building and evaluating complex information processing models of the human brain ([119, 130, 214]).
@@
Despite the remarkable advances in AI tools and the new opportunities that ANNs present for MEG data analysis and modeling, it is critical to avoid the allure of AI hype and the rush to employ complex AI algorithms solely for their novelty. Many research questions may still be best addressed using standard approaches that are well-established within the field and which often come with greater result interpretability. Moreover, the extensive range of parameters and hyperparameters involved in ANNs, coupled with the expertise required for their proper implementation, can sometimes introduce errors or yield misleading outcomes.
@@
Looking ahead, the future of this emerging research appears ripe with promising developments. Two particularly exciting prospects stand out. First, advances in AI interpretability and visualization tools are poised to significantly enhance the application of ANNs in MEG research, making these complex models more accessible and their findings more actionable. Second, the burgeoning adoption of AI foundation models across various disciplines, including neuroscience, is set to play a transformative role in both basic and clinical MEG applications. The scientific community’s growing enthusiasm for developing foundation models for brain imaging modalities ([199, 215]) and time series data ([216]) suggests that neuroscience will not be far behind ([162, 114]). Several studies have successfully applied ANNs to multimodal neuroimaging data, including MEG-EEG integration. For example, frameworks have been developed to process and fuse MEG and EEG signals for enhanced decoding performance and artifact removal using deep neural architectures ([217, 218, 219]). These approaches highlight the potential of ANNs to handle the complementary strengths of different modalities and suggest promising avenues for future multimodal brain decoding pipelines. Integrating such modalities could revolutionize our understanding of the brain and its disorders.
Acknowledgements
Conflict of interest
The authors declare that they have no conflicts of interest relevant to this work.
Funding
I.R. received support from the Canada CIFAR AI chair program and the Canada excellence research chairs program. K.J. was supported by Canada research chairs program funding (950-232368). V.H. was supported through the Canadian Institutes of health tesearch project grant (166197). H.A. was funded through the faculty of medicine at the university of Montreal’s merit scholarship. A.D. was funded through computational neuroscience and cognitive neuroimaging lab (CoCoLab) and Montreal institute of learning algorithms (MILA) Scholarships.